# Irregular Tensor Low-Rank Representation for Hyperspectral Image Representation

### v1

@@ IV-D Ablation Study
# Origin LSSTRPCA $\mathrm{S^{3}LRR}$ LPGTRPCA TensorSSA Proposed
1 97.19 ± 1.88 97.73 ± 2.07 95.74 ± 2.45 98.42 ± 0.88 98.35 ± 2.07 97.78 ± 3.21
2 99.15 ± 0.54 98.11 ± 1.72 96.67 ± 2.34 98.15 ± 1.22 99.49 ± 0.67 99.62 ± 0.48
3 97.22 ± 1.88 91.16 ± 4.51 96.85 ± 2.89 97.40 ± 2.20 97.57 ± 2.11 99.94 ± 0.12
4 99.07 ± 0.85 96.01 ± 4.33 90.37 ± 6.48 93.52 ± 3.00 98.88 ± 0.53 95.58 ± 2.90
5 96.96 ± 1.24 93.27 ± 4.24 91.28 ± 3.88 95.11 ± 2.26 96.73 ± 1.60 97.85 ± 0.95
6 99.21 ± 0.66 99.67 ± 0.36 98.73 ± 0.81 99.04 ± 0.41 99.78 ± 0.09 99.89 ± 0.11
7 99.41 ± 0.23 98.43 ± 0.79 97.55 ± 1.49 97.50 ± 1.35 99.32 ± 0.49 99.51 ± 0.44
8 82.33 ± 1.90 96.12 ± 1.73 97.21 ± 1.13 95.06 ± 1.79 94.04 ± 1.87 97.74 ± 1.38
9 98.90 ± 0.93 98.99 ± 0.88 98.86 ± 1.60 99.33 ± 0.84 99.46 ± 0.73 99.51 ± 0.38
10 88.87 ± 4.45 91.96 ± 2.42 91.16 ± 2.99 92.64 ± 3.47 93.70 ± 2.79 95.39 ± 2.10
11 91.38 ± 4.33 90.37 ± 6.62 84.91 ± 4.74 89.42 ± 7.22 95.13 ± 4.96 96.67 ± 4.36
12 98.99 ± 1.17 97.70 ± 4.06 94.18 ± 4.01 94.23 ± 3.63 99.93 ± 0.06 99.42 ± 0.79
13 95.23 ± 4.98 94.44 ± 3.95 88.75 ± 7.22 90.06 ± 6.64 97.49 ± 1.22 95.91 ± 2.87
14 91.65 ± 4.22 93.30 ± 3.22 90.10 ± 5.53 87.61 ± 7.65 93.27 ± 5.73 96.39 ± 3.32
15 67.96 ± 2.58 95.88 ± 1.90 97.16 ± 1.73 95.75 ± 1.37 94.50 ± 2.08 97.44 ± 1.56
16 95.82 ± 3.38 96.87 ± 2.93 97.62 ± 2.53 97.26 ± 1.64 95.10 ± 5.80 98.09 ± 1.06
OA 90.09 ± 0.40 96.33 ± 0.46 95.96 ± 0.55 96.03 ± 0.63 96.70 ± 0.51 98.19 ± 0.41
AA 93.71 ± 8.34 95.63 ± 2.90 94.20 ± 4.19 95.03 ± 3.59 97.05 ± 2.41 97.92 ± 1.60
$\kappa$ 88.96 ± 0.45 95.92 ± 0.51 95.51 ± 0.61 95.58 ± 0.70 96.33 ± 0.57 97.98 ± 0.45
TABLE XIII: CLASSIFICATION PERFORMANCE OF TESTED METHODS ON PAVIA UNIVERSITY WITH 0.5% TRAINING SAMPLES. OPTIMAL VALUES ARE DENOTED IN BOLD, AND THE SECOND-BEST VALUES ARE UNDERLINED.
# Origin LSSTRPCA $\mathrm{S^{3}LRR}$ LPGTRPCA TensorSSA Proposed
1 85.66 ± 3.22 85.34 ± 3.35 88.61 ± 2.98 85.27 ± 3.50 90.97 ± 3.08 86.65 ± 2.90
2 93.90 ± 2.40 96.52 ± 1.13 95.07 ± 1.80 95.31 ± 1.99 96.92 ± 1.62 99.25 ± 0.55
3 63.30 ± 10.42 64.03 ± 8.24 63.24 ± 9.24 67.87 ± 5.73 82.31 ± 6.98 75.72 ± 7.40
4 81.76 ± 5.69 90.48 ± 4.11 82.06 ± 8.29 86.98 ± 4.95 87.05 ± 3.55 90.87 ± 2.64
5 94.79 ± 10.24 98.65 ± 0.99 99.19 ± 0.72 97.88 ± 3.42 95.73 ± 6.97 97.94 ± 4.70
6 74.12 ± 5.43 80.02 ± 5.99 85.74 ± 6.67 74.61 ± 8.31 85.76 ± 4.39 95.26 ± 2.63
7 68.70 ± 10.47 74.03 ± 8.77 80.90 ± 10.05 72.06 ± 10.17 88.51 ± 6.27 82.44 ± 8.04
8 77.05 ± 6.35 78.41 ± 5.83 74.77 ± 7.08 80.21 ± 5.67 83.60 ± 4.82 87.74 ± 3.87
9 99.72 ± 0.13 99.68 ± 0.15 99.50 ± 0.29 99.76 ± 0.13 98.63 ± 0.59 99.70 ± 0.27
OA 85.85 ± 1.41 88.70 ± 1.20 88.52 ± 1.51 87.53 ± 1.15 91.85 ± 1.24 93.53 ± 0.73
AA 82.11 ± 12.48 85.24 ± 12.21 85.45 ± 11.88 84.44 ± 11.63 89.94 ± 5.97 90.62 ± 8.26
$\kappa$ 81.12 ± 1.83 84.95 ± 1.61 84.72 ± 2.01 83.36 ± 1.55 89.15 ± 1.63 91.40 ± 0.97
@@
TABLE XIV: CLASSIFICATION PERFORMANCE OF TESTED METHODS ON WHU-HI-LONGKOU WITH 1% TRAINING SAMPLES. OPTIMAL VALUES ARE DENOTED IN BOLD, AND THE SECOND-BEST VALUES ARE UNDERLINED.
@@
# Origin LSSTRPCA $\mathrm{S^{3}LRR}$ LPGTRPCA TensorSSA Proposed
1 88.38 ± 4.13 91.74 ± 5.37 92.58 ± 3.81 94.96 ± 3.27 96.78 ± 1.54 99.08 ± 0.75
2 87.88 ± 3.32 95.29 ± 2.05 91.94 ± 2.67 94.61 ± 2.02 95.93 ± 2.67 98.09 ± 1.49
3 73.27 ± 6.44 87.20 ± 5.42 84.90 ± 4.43 91.53 ± 4.23 97.40 ± 3.12 95.16 ± 1.75
4 85.51 ± 2.83 91.09 ± 2.40 89.62 ± 2.32 91.09 ± 2.61 96.37 ± 1.43 95.86 ± 1.79
5 71.11 ± 5.00 73.98 ± 5.01 76.98 ± 6.18 69.09 ± 8.42 90.74 ± 3.90 92.64 ± 2.82
6 87.43 ± 7.51 95.85 ± 4.51 93.11 ± 6.46 96.28 ± 3.57 93.64 ± 5.41 99.98 ± 0.07
7 99.98 ± 0.01 99.95 ± 0.04 $9$ 0.01 99.88 ± 0.13 99.87 ± 0.08 99.94 ± 0.04
8 79.61 ± 6.65 80.36 ± 4.41 84.19 ± 4.26 84.88 ± 6.23 84.62 ± 5.82 90.23 ± 3.39
9 55.93 ± 5.65 54.32 ± 6.80 58.43 ± 7.29 61.91 ± 5.46 77.07 ± 8.72 80.33 ± 8.28
OA 91.23 ± 0.66 93.54 ± 0.58 93.44 ± 0.64 94.11 ± 0.56 96.53 ± 0.55 97.36 ± 0.46
AA 81.01 ± 12.81 85.53 ± 14.21 85.75 ± 12.16 87.14 ± 13.06 92.49 ± 7.31 94.59 ± 6.31
$\kappa$ 86.96 ± 0.97 90.39 ± 0.86 90.25 ± 0.95 91.24 ± 0.83 94.85 ± 0.81 96.08 ± 0.68
IV-E Comparison with State-of-the-Art Methods

### v3

@@ IV-A Datasets
# Train Test Origin SpeFormer [46] DSNet [43] HyperDID [59] LSSTRPCA [23] $\mathrm{S^{3}LRR}$ [12] LPGTRPCA [24] TensorSSA [60] Proposed DSNet+Ours
1 5 41 63.75 ± 14.75 40.00 ± 14.00 62.44 ± 20.15 8.78 ± 17.56 85.50 ± 9.23 85.63 ± 12.35 91.25 ± 9.41 92.75 ± 5.84 94.00 ± 1.37 $\bullet$90.73 ± 10.39
2 143 1285 79.19 ± 2.49 88.95 ± 3.69 82.51 ± 10.93 94.23 ± 2.56 93.93 ± 2.09 97.25 ± 1.20 97.11 ± 1.22 96.69 ± 1.12 98.72 ± 0.77 $\bullet$98.54 ± 0.81
3 83 747 71.03 ± 3.02 88.15 ± 4.19 88.78 ± 4.72 93.04 ± 2.84 92.73 ± 3.16 95.32 ± 2.19 96.39 ± 1.61 96.92 ± 1.64 96.86 ± 1.38 $\bullet$99.20 ± 1.10
4 24 213 55.00 ± 6.32 77.80 ± 6.30 68.26 ± 12.19 83.76 ± 8.62 91.22 ± 5.66 94.23 ± 3.74 97.58 ± 2.10 96.92 ± 3.01 94.46 ± 2.88 $\bullet$97.84 ± 1.56
5 49 434 90.18 ± 2.33 92.97 ± 2.98 88.80 ± 4.60 96.91 ± 1.43 92.51 ± 2.61 94.99 ± 2.09 96.10 ± 2.44 95.02 ± 2.48 98.38 ± 1.02 $\bullet$99.08 ± 0.73
6 73 657 95.60 ± 1.50 97.46 ± 2.26 99.33 ± 0.72 99.36 ± 0.51 99.32 ± 0.65 99.50 ± 0.46 99.02 ± 0.77 99.18 ± 0.67 99.54 ± 1.02 $\bullet$99.88 ± 0.18
7 3 25 75.80 ± 12.34 39.23 ± 18.90 63.20 ± 21.97 7.20 ± 14.40 71.00 ± 18.44 70.80 ± 18.72 96.20 ± 6.15 91.80 ± 7.73 88.00 ± 12.00 $\bullet$64.80 ± 17.42
8 48 430 96.84 ± 1.81 99.12 ± 0.78 99.07 ± 1.03 99.81 ± 0.37 98.82 ± 1.79 99.69 ± 0.75 99.74 ± 0.49 99.24 ± 0.96 100.00 ± 0.00 $\bullet$99.86 ± 0.28
9 2 18 37.22 ± 13.50 40.00 ± 25.56 43.33 ± 23.41 0.00 ± 0.00 56.67 ± 17.44 71.67 ± 15.39 68.06 ± 22.43 96.94 ± 7.09 82.22 ± 18.17 $\bullet$66.67 ± 22.22
10 98 874 74.30 ± 2.69 90.11 ± 4.29 93.00 ± 3.28 92.01 ± 3.06 91.21 ± 2.21 94.43 ± 1.81 95.93 ± 1.57 96.39 ± 1.75 99.11 ± 0.83 $\bullet$99.57 ± 0.29
11 246 2209 79.41 ± 2.44 94.32 ± 2.14 96.04 ± 2.21 95.27 ± 2.06 95.47 ± 1.15 97.95 ± 0.82 97.69 ± 0.64 98.75 ± 0.48 99.38 ± 0.40 $\bullet$99.33 ± 0.47
12 60 533 70.37 ± 3.26 81.65 ± 4.74 86.04 ± 5.76 88.93 ± 1.88 93.40 ± 2.35 93.39 ± 2.52 95.08 ± 2.20 94.60 ± 2.06 98.46 ± 0.93 $\bullet$97.71 ± 1.63
13 21 184 96.42 ± 2.37 98.86 ± 1.42 99.13 ± 1.01 99.89 ± 0.22 96.64 ± 3.30 98.44 ± 2.49 99.40 ± 1.03 98.63 ± 1.18 100.00 ± 0.00 $\bullet$99.89 ± 0.22
14 127 1138 93.00 ± 1.85 96.62 ± 2.11 96.40 ± 1.43 97.35 ± 1.29 98.39 ± 0.94 99.19 ± 0.58 99.40 ± 0.39 99.39 ± 0.59 99.95 ± 0.12 $\bullet$99.93 ± 0.10
15 39 347 52.80 ± 6.37 77.30 ± 7.36 68.88 ± 19.29 79.77 ± 10.11 98.85 ± 1.79 98.49 ± 1.15 98.71 ± 1.56 96.72 ± 2.53 99.60 ± 0.26 $\bullet$99.08 ± 0.74
16 10 83 89.58 ± 4.46 87.62 ± 7.57 99.28 ± 1.45 98.31 ± 2.81 91.02 ± 6.21 90.00 ± 6.26 98.19 ± 3.14 96.27 ± 4.69 90.36 ± 2.69 $\circ$94.70 ± 3.54
OA 80.54 ± 0.83 91.10 ± 0.90 90.86 ± 1.64 93.57 ± 0.42 94.99 ± 0.50 96.94 ± 0.40 97.50 ± 0.34 97.62 ± 0.32 98.81 ± 0.21 $\bullet$98.99 ± 0.32
AA 76.28 ± 17.56 80.64 ± 3.01 83.41 ± 2.78 77.16 ± 1.52 90.42 ± 11.32 92.56 ± 9.12 95.36 ± 7.58 96.64 ± 2.24 96.19 ± 5.20 $\bullet$94.17 ± 2.18
$\kappa$ 77.80 ± 0.93 89.83 ± 1.04 89.55 ± 1.88 92.66 ± 0.49 94.28 ± 0.58 96.51 ± 0.46 97.14 ± 0.39 97.28 ± 0.37 98.64 ± 0.24 $\bullet$98.85 ± 0.37
Time (s) 33 262 302 127 659 570 621 59 1102 -
@@
TABLE IV: Classification Performance of Comparison Methods on Salinas with 1.5% Training Samples. Optimal Values Are Denoted in Bold, and the Second-Best Values Are Underlined. $\bullet$/$\circ$ Indicates Whether the Performance of DSNet on Data Processed by Our Method Is Superior/Inferior to That on the Original Data
Deep learning methods Tensor decomposition methods Ours
@@
# Train Test Origin SpeFormer [46] DSNet [43] HyperDID [59] LSSTRPCA [23] $\mathrm{S^{3}LRR}$ [12] LPGTRPCA [24] TensorSSA [60] Proposed DSNet+Ours
1 31 1978 98.69 ± 1.34 93.33 ± 9.55 95.28 ± 5.62 97.14 ± 2.93 97.89 ± 0.77 96.36 ± 3.30 98.88 ± 0.40 99.24 ± 1.23 99.86 ± 0.24 $\bullet$99.97 ± 0.06
2 56 3670 99.40 ± 0.47 98.85 ± 0.75 99.97 ± 0.20 99.18 ± 1.13 98.32 ± 0.92 97.87 ± 1.97 98.58 ± 1.16 99.34 ± 0.83 99.95 ± 0.08 $\circ$99.91 ± 0.19
3 30 1946 98.00 ± 1.57 79.01 ± 16.06 98.81 ± 1.31 98.61 ± 0.68 94.55 ± 3.22 96.61 ± 2.80 95.70 ± 4.74 98.84 ± 0.49 99.87 ± 0.25 $\bullet$99.98 ± 0.04
4 21 1373 99.40 ± 0.29 98.30 ± 0.67 98.79 ± 1.52 99.40 ± 0.52 97.92 ± 1.07 92.37 ± 4.01 94.80 ± 5.07 98.99 ± 0.29 96.37 ± 1.08 $\circ$92.88 ± 7.86
5 41 2637 96.87 ± 1.40 98.19 ± 1.51 99.64 ± 0.29 98.74 ± 1.25 95.22 ± 2.12 95.06 ± 2.84 96.36 ± 1.30 96.65 ± 1.47 98.23 ± 0.52 $\circ$98.46 ± 0.95
6 60 3899 99.38 ± 0.42 99.79 ± 0.37 100.00 ± 0.00 99.93 ± 0.13 99.84 ± 0.14 98.96 ± 0.83 99.63 ± 0.30 99.87 ± 0.09 99.94 ± 0.01 $\circ$99.90 ± 0.14
7 54 3525 99.52 ± 0.16 97.49 ± 1.55 99.43 ± 0.18 99.41 ± 0.38 98.59 ± 1.06 98.13 ± 0.93 97.65 ± 0.85 99.60 ± 0.26 99.80 ± 0.06 $\bullet$99.81 ± 0.26
8 170 11101 82.50 ± 1.19 83.07 ± 8.43 89.71 ± 4.53 90.75 ± 2.90 98.50 ± 0.44 98.23 ± 0.38 97.55 ± 1.20 95.00 ± 0.68 98.22 ± 1.27 $\bullet$99.55 ± 0.18
9 94 6109 99.09 ± 0.76 98.77 ± 0.59 99.68 ± 0.52 99.74 ± 0.26 98.90 ± 0.84 98.35 ± 1.45 99.77 ± 0.25 99.50 ± 0.43 99.87 ± 0.14 $\bullet$99.76 ± 0.32
10 50 3228 93.18 ± 1.28 92.03 ± 3.08 96.38 ± 1.05 94.59 ± 1.22 94.31 ± 1.42 95.57 ± 0.96 97.68 ± 1.07 95.70 ± 1.60 95.84 ± 1.11 $\bullet$98.46 ± 1.82
11 17 1051 93.53 ± 4.81 88.54 ± 5.19 97.37 ± 1.73 94.39 ± 5.08 94.23 ± 2.54 89.46 ± 8.53 91.95 ± 2.74 97.94 ± 1.93 98.71 ± 1.95 $\circ$96.82 ± 5.66
12 29 1898 98.89 ± 1.69 98.56 ± 2.29 100.00 ± 0.00 98.68 ± 2.63 98.96 ± 0.69 95.34 ± 6.37 96.72 ± 0.48 99.97 ± 0.05 99.60 ± 0.41 $\circ$99.40 ± 0.84
13 14 902 98.12 ± 0.92 98.43 ± 1.81 99.87 ± 0.21 99.84 ± 0.09 94.30 ± 3.38 84.17 ± 10.19 91.82 ± 1.96 97.98 ± 1.50 97.69 ± 1.42 $\circ$94.41 ± 5.46
14 17 1053 91.02 ± 2.19 95.21 ± 2.03 98.25 ± 0.70 94.85 ± 1.51 93.66 ± 2.67 93.90 ± 2.64 93.62 ± 8.35 95.78 ± 2.69 97.17 ± 0.87 $\circ$98.12 ± 1.79
15 110 7158 71.08 ± 3.27 75.95 ± 7.91 72.94 ± 12.49 75.14 ± 3.97 97.72 ± 0.62 98.29 ± 0.90 96.87 ± 1.35 96.28 ± 1.66 98.61 ± 0.46 $\bullet$99.42 ± 0.21
16 28 1779 97.31 ± 0.94 92.46 ± 2.91 98.85 ± 0.62 97.64 ± 1.36 97.97 ± 1.12 97.07 ± 3.58 99.07 ± 0.54 98.23 ± 0.49 98.23 ± 0.16 $\bullet$99.33 ± 0.66
91.08 ± 0.43 90.56 ± 1.61 93.53 ± 1.23 93.70 ± 0.48 97.68 ± 0.28 97.04 ± 0.47 97.54 ± 0.28 97.58 ± 0.33 98.77 ± 0.32 $\bullet$99.18 ± 0.16
AA 94.75 ± 7.75 93.00 ± 1.81 96.56 ± 0.76 96.13 ± 0.29 96.93 ± 2.12 95.36 ± 3.91 96.66 ± 2.52 98.06 ± 1.66 98.62 ± 1.33 $\bullet$98.51 ± 0.37
$\kappa$ 90.06 ± 0.48 89.50 ± 1.79 92.79 ± 1.39 92.98 ± 0.53 97.42 ± 0.32 96.70 ± 0.53 97.26 ± 0.32 97.31 ± 0.37 98.63 ± 0.35 $\bullet$99.09 ± 0.18
Time (s) 39 1281 898 271 1452 1892 1506 314 6015 -
TABLE V: Classification Performance of Comparison Methods on Pavia University with 0.5% Training Samples. Optimal Values Are Denoted in Bold, and the Second-Best Values Are Underlined. $\bullet$/$\circ$ Indicates Whether the Performance of DSNet on Data Processed by Our Method Is Superior/Inferior to That on the Original Data
Deep learning methods Tensor decomposition methods Ours
# Train Test Origin SpeFormer [46] DSNet [43] HyperDID [59] LSSTRPCA [23] $\mathrm{S^{3}LRR}$ [12] LPGTRPCA [24] TensorSSA [60] Proposed DSNet+Ours
1 34 6597 85.66 ± 3.22 73.15 ± 6.86 90.64 ± 5.17 94.00 ± 0.88 85.34 ± 3.35 88.61 ± 2.98 85.27 ± 3.50 90.97 ± 3.08 86.65 ± 2.90 $\bullet$92.25 ± 7.13
2 94 18555 93.90 ± 2.40 95.77 ± 3.57 99.72 ± 0.14 98.08 ± 0.90 96.52 ± 1.13 95.07 ± 1.80 95.31 ± 1.99 96.92 ± 1.62 99.25 ± 0.55 $\circ$99.24 ± 0.96
3 11 2088 63.30 ± 10.42 48.52 ± 11.24 66.88 ± 20.78 43.74 ± 12.03 64.03 ± 8.24 63.24 ± 9.24 67.87 ± 5.73 82.31 ± 6.98 75.72 ± 7.40 $\circ$62.80 ± 17.66
4 16 3048 81.76 ± 5.69 85.41 ± 4.41 89.97 ± 3.28 90.77 ± 5.46 90.48 ± 4.11 82.06 ± 8.29 86.98 ± 4.95 87.05 ± 3.55 90.87 ± 2.64 $\bullet$93.76 ± 1.29
5 7 1338 94.79 ± 10.24 99.40 ± 0.75 96.82 ± 3.52 98.89 ± 1.25 98.65 ± 0.99 99.19 ± 0.72 97.88 ± 3.42 95.73 ± 6.97 97.94 ± 4.70 $\bullet$99.48 ± 0.43
6 26 5003 74.12 ± 5.43 42.06 ± 12.26 84.27 ± 4.34 84.22 ± 3.57 80.02 ± 5.99 85.74 ± 6.67 74.61 ± 8.31 85.76 ± 4.39 95.26 ± 2.63 $\bullet$98.14 ± 1.33
7 7 1323 68.70 ± 10.47 62.12 ± 6.83 78.82 ± 17.91 70.04 ± 3.58 74.03 ± 8.77 80.90 ± 10.05 72.06 ± 10.17 88.51 ± 6.27 82.44 ± 8.04 $\bullet$82.46 ± 13.68
8 19 3663 77.05 ± 6.35 74.48 ± 8.66 90.37 ± 4.69 84.64 ± 8.10 78.41 ± 5.83 74.77 ± 7.08 80.21 ± 5.67 83.60 ± 4.82 87.74 ± 3.87 $\bullet$93.75 ± 6.10
9 5 942 99.72 ± 0.13 92.21 ± 3.53 96.82 ± 2.65 99.15 ± 0.97 99.68 ± 0.15 99.50 ± 0.29 99.76 ± 0.13 98.63 ± 0.59 99.70 ± 0.27 $\bullet$99.00 ± 1.16
OA 85.85 ± 1.41 80.05 ± 1.69 92.58 ± 1.33 90.65 ± 0.47 88.70 ± 1.20 88.52 ± 1.51 87.53 ± 1.15 91.85 ± 1.24 93.53 ± 0.73 $\bullet$94.85 ± 1.73
AA 82.11 ± 12.48 74.79 ± 1.86 88.26 ± 2.20 84.84 ± 0.78 85.24 ± 12.21 85.45 ± 11.88 84.44 ± 11.63 89.94 ± 5.97 90.62 ± 8.26 $\bullet$91.21 ± 3.59
$\kappa$ 81.12 ± 1.83 72.95 ± 2.36 90.07 ± 1.78 87.50 ± 0.59 84.95 ± 1.61 84.72 ± 2.01 83.36 ± 1.55 89.15 ± 1.63 91.40 ± 0.97 $\bullet$93.17 ± 2.30
Time (s) 11 969 697 246 1895 2449 1392 324 3751 -
TABLE VI: Classification Performance of Comparison Methods on WHU-Hi-LongKou with 1% Training Samples. Optimal Values Are Denoted in Bold, and the Second-Best Values Are Underlined. $\bullet$/$\circ$ Indicates Whether the Performance of DSNet on Data Processed by Our Method Is Superior/Inferior to That on the Original Data
Deep learning methods Tensor decomposition methods Ours
# Train Test Origin SpeFormer [46] DSNet [43] HyperDID [59] LSSTRPCA [23] $\mathrm{S^{3}LRR}$ [12] LPGTRPCA [24] TensorSSA [60] Proposed DSNet+Ours
1 15 1396 88.38 ± 4.13 92.39 ± 2.60 96.88 ± 1.70 97.03 ± 1.58 91.74 ± 5.37 92.58 ± 3.81 94.96 ± 3.27 96.78 ± 1.54 99.08 ± 0.75 $\bullet$99.70 ± 0.53
2 36 3532 87.88 ± 3.32 88.78 ± 3.57 96.70 ± 1.18 96.19 ± 2.73 95.29 ± 2.05 91.94 ± 2.67 94.61 ± 2.02 95.93 ± 2.67 98.09 ± 1.49 $\bullet$99.39 ± 0.28
3 13 1216 73.27 ± 6.44 68.87 ± 23.91 96.38 ± 1.58 55.44 ± 29.45 87.20 ± 5.42 84.90 ± 4.43 91.53 ± 4.23 97.40 ± 3.12 95.16 ± 1.75 $\bullet$98.95 ± 1.30
4 59 5773 85.51 ± 2.83 66.73 ± 8.56 88.79 ± 3.38 90.80 ± 4.29 91.09 ± 2.40 89.62 ± 2.32 91.09 ± 2.61 96.37 ± 1.43 95.86 ± 1.79 $\bullet$96.55 ± 1.67
5 14 1324 71.11 ± 5.00 74.15 ± 5.14 77.61 ± 4.17 48.69 ± 30.34 73.98 ± 5.01 76.98 ± 6.18 69.09 ± 8.42 90.74 ± 3.90 92.64 ± 2.82 $\bullet$95.47 ± 3.19
6 7 665 87.43 ± 7.51 93.80 ± 5.88 99.28 ± 0.93 97.02 ± 5.08 95.85 ± 4.51 93.11 ± 6.46 96.28 ± 3.57 93.64 ± 5.41 99.98 ± 0.07 $\bullet$99.97 ± 0.06
7 207 20403 99.98 ± 0.01 99.95 ± 0.07 99.83 ± 0.06 $9$ 0.01 99.95 ± 0.04 $9$ 0.01 99.88 ± 0.13 99.87 ± 0.08 99.94 ± 0.04 $\circ$99.71 ± 0.17
8 25 2428 79.61 ± 6.65 80.71 ± 10.24 88.36 ± 4.55 90.74 ± 6.02 80.36 ± 4.41 84.19 ± 4.26 84.88 ± 6.23 84.62 ± 5.82 90.23 ± 3.39 $\bullet$93.52 ± 6.10
9 15 1435 55.93 ± 5.65 80.47 ± 6.92 75.16 ± 6.13 43.48 ± 21.76 54.32 ± 6.80 58.43 ± 7.29 61.91 ± 5.46 77.07 ± 8.72 80.33 ± 8.28 $\bullet$84.68 ± 6.56
OA 91.23 ± 0.66 89.66 ± 1.26 95.21 ± 0.43 92.18 ± 1.38 93.54 ± 0.58 93.44 ± 0.64 94.11 ± 0.56 96.53 ± 0.55 97.36 ± 0.46 $\bullet$98.07 ± 0.29
AA 81.01 ± 12.81 82.87 ± 3.32 91.00 ± 0.69 79.93 ± 4.79 85.53 ± 14.21 85.75 ± 12.16 87.14 ± 13.06 92.49 ± 7.31 94.59 ± 6.31 $\bullet$96.44 ± 0.54
$\kappa$ 86.96 ± 0.97 84.72 ± 1.86 92.90 ± 0.63 88.31 ± 2.08 90.39 ± 0.86 90.25 ± 0.95 91.24 ± 0.83 94.85 ± 0.81 96.08 ± 0.68 $\bullet$97.14 ± 0.43
Time (s) 15 882 669 199 791 1628 815 127 1311 -
IV-B Experiment Setup and Compared Methods
To evaluate the representation ability, we performed the classification task on the learned representation. Specifically, a typical SVM classifier equipped with an RBF kernel was implemented as the classifier, and three metrics, including the overall classification accuracy (OA), the average class classification accuracy (AA), and the kappa coefficient ($\kappa$), were used to measure the classification accuracy. We repeated the classifier 5 times to obtain the average of OA, AA, and $\kappa$ across all compared methods.
We compared the classification performance of our method with seven state-of-the-art methods, including four tensor-based methods, i.e., LSSTRPCA [23], $\mathrm{S^{3}LRR}$ [12], LPGTRPCA [24], TensorSSA [60] and three deep learning methods, i.e., SpeFormer [46], DSNet [43], HyperDID [59]. For a fair comparison, we used the codes provided by their inventors, and the parameter settings of these methods were tuned to optimal ones. Table II presents the parameter settings for the proposed method. The details of the compared methods are listed as follows:
  1. 1.
    LSSTRPCA [23]: is a lateral-slice sparse tensor RPCA with a tensor $l_{2,1}$ norm for sparse component to gross errors or outliers.
  2. 2.
    $\mathrm{S^{3}LRR}$ [12]: is a model exploring the low-rank property spectral and spatial domain simultaneously in the 2-D matrix domain.
  3. 3.
    LPGTRPCA [24]: is a tensor RPCA model with a locality-preserving graph and frontal slice sparsity.
  4. 4.
    TensorSSA [60]: is a method of 3D tensor Singular Spectrum Analysis (SSA) that extracts the low-rank features of HSI by decomposing the trajectory tensor and reconstructing it with low-rank approximation.
  5. 5.
    SpeFormer [46]: is a transformer-based model for the HSI classification task, leveraging group-wise spectral embeddings and cross-layer adaptive fusion to improve the capture of spectral and spatial features.
  6. 6.
    DSNet [43]: is a dual-branch subpixel-guided network that integrates a deep autoencoder unmixing architecture to enhance hyperspectral image classification by fusing subpixel and pixel-level features.
  7. 7.
    HyperDID [59]: is a hyperspectral intrinsic image decomposition framework incorporating a deep feature embedding to separate environment-related and category-related features for improved hyperspectral image classification.
All the tensor-based methods were conducted on a Windows 10 server equipped with two Intel Xeon Gold 6248R CPUs. The deep-learning methods were carried out on a Linux server with an AMD EPYC 7642 CPU and an NVIDIA RTX 4090 GPU.
Figure 8: Influence of the two hyper-parameters ($\alpha$ and $\beta$) of our model with respect to OA (%) on four datasets.
IV-C Comparison with State-of-the-Art Methods